OpenStudy (anonymous):

simplify : underroot5-1 / underroot5+1

jhonyy9 (jhonyy9):

underroot wann signing squarroot ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so than will be (sqrt5-1)2 squared)/4

jhonyy9 (jhonyy9):

multiply numerator and denominator by sqrt5-1

jhonyy9 (jhonyy9):

using formula a2 -b2 = (a-b)(a+b)

OpenStudy (anonymous):

its \[\sqrt{5}-1\div \sqrt{5}+1\]

jhonyy9 (jhonyy9):

yes i understand and those is the way wath you need to do

jhonyy9 (jhonyy9):

and this will be the resulted answer

OpenStudy (anonymous):

so what isthe final answer ?

OpenStudy (king):

(2-sqrt5)/2